# Shape Formation and Locomotion with Joint Movements in the Amoebot Model

### v2

@@ 5.4 Transportation
Figure 21: Transportation of objects by the millipede structure. The green rhombi indicate the millipede structure. The gray parallelogram indicates the object.
@@ 6 Conclusion and Future Work
In this paper, we have formalized the joint movement extension that were proposed by Feldmann et al. [23]. We have constructed meta-modules of rhombical and hexagonal shape that are able to perform various movement primitives. This allows us to simulate shape formation algorithms of various MRSs. However, our meta-modules are more flexible, e.g., we can move a hexagonal meta-module through two others (see Appendix A). Such new movement primitives may lead to faster shape formation algorithms, e.g., sublinear solutions for hexagonal meta-modules or even arbitrary amoebot structures. Furthermore, we have presented three amoebot structures capable of moving along an even surface. In future work, movement on uneven surfaces can be considered.
@@ Appendix A Moving Meta-Modules through Others
In this section, we show how to move a hexagonal meta-module $H_{1}$ through two other hexagonal meta-modules $H_{2}$ and $H_{3}$ (see Figure 22). We first form six rhombical meta-modules from $H_{1}$ and the amoebots of $H_{2}$ and $H_{3}$ that are located between the old and new position of $H_{1}$. For that, we shift the amoebots along the red lines. Using the realignment primitive and the slide primitive for rhombical meta-modules, we shift the six rhombical meta-modules into the new position of $H_{1}$. Finally, we restore $H_{2}$ and $H_{3}$ by shifting the amoebots along the red lines. Note that we have omitted some necessary reorientations.
@@
Figure 22: Moving a hexagonal meta-module through two other hexagonal meta-modules. For the sake of clarity, we have omitted the bonds in these figures. In the first and last step, the amoebots move along the red lines.

### v3

@@ 5.4 Transportation
Figure 34: Transportation of objects by the millipede structure. The arrow indicates two rounds. The green rhombi indicate the millipede structure. The gray parallelogram indicates the object.
@@ 6 Conclusion and Future Work
In this paper, we have formalized the joint movement extension that were proposed by Feldmann et al. [23]. We have constructed meta-modules of rhombical and hexagonal shape that are able to perform various movement primitives. This allows us to simulate reconfiguration algorithms of various MRSs. However, our meta-modules are more flexible, e.g., we can move a hexagonal meta-module through two others (see Appendix A).
Such new movement primitives may lead to faster reconfiguration algorithms, e.g., sublinear solutions for hexagonal meta-modules or even arbitrary amoebot structures. Furthermore, we have presented three amoebot structures capable of moving along an even surface. In future work, movement on uneven surfaces can be considered.
@@ Appendix A Moving Meta-Modules through Others
In this section, we show how to move a hexagonal meta-module $H_{1}$ through two other hexagonal meta-modules $H_{2}$ and $H_{3}$ (see Figure 35). We first form six rhombical meta-modules from $H_{1}$ and the amoebots of $H_{2}$ and $H_{3}$ that are located between the old and new position of $H_{1}$. For that, we shift the amoebots along the red lines. In order to shift the amoebots along the red lines, we first contract all amoebots on the line and then expand all amoebots on the line into the other direction. Using the realignment primitive and the slide primitive for rhombical meta-modules, we shift the six rhombical meta-modules into the new position of $H_{1}$. Finally, we restore $H_{2}$ and $H_{3}$ by shifting the amoebots along the red lines. Note that we have omitted some necessary reorientations.
@@
Figure 35: Moving a hexagonal meta-module through two other hexagonal meta-modules. The arrows indicate multiple rounds, respectively. For the sake of clarity, we have omitted the bonds in these figures. In the first and last step, the amoebots move along the red lines.